OpenStudy (anonymous):

When solving the system x + 2y = 9 and x - y = 6, which expression could be substituted for x in the second equation?

OpenStudy (anonymous):

A. 9 - x B. y - 9/2 C. 9 - 2y D. 1/2 y + 9

OpenStudy (anonymous):

x - y = 6 x = 6+y x + 2y = 9 x = 9-2y

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

nom. questions. should be questioned! lol. I'm kidding. do you understand why? understanding means you can do 100's of these whereas posting 100's of these is hard to respond to o_o